OpenStudy (anonymous):

What is the domain of f(x)=1/(x+1)-4

OpenStudy (anonymous):

Would it be all numbers except 3? @Eyad

OpenStudy (anonymous):

Thats right :)

OpenStudy (anonymous):

Thanks so much! =)

OpenStudy (anonymous):

you 're welcome so much ;)

jimthompson5910 (jim_thompson5910):

no, 3 is a valid x value to plug in

jimthompson5910 (jim_thompson5910):

1/(x+1)+4 1/(3+1)+4 1/4+4 17/4 So 1/(x+1)+4 = 17/4 when x = 3

OpenStudy (anonymous):

@jim_thompson5910 : its (x+1)-4 , not (x+1)+4

jimthompson5910 (jim_thompson5910):

oh right, my bad, but it still means that the domain of \[\Large f(x)=\frac{1}{x+1}-4\] isn't every real number but 3

OpenStudy (anonymous):

the denominator is :[ (x+1)-4 ] x-3=0 , x=3 .

jimthompson5910 (jim_thompson5910):

since 3 is still a valid input

jimthompson5910 (jim_thompson5910):

are you sure?

jimthompson5910 (jim_thompson5910):

I'm thinking that the denominator is just x+1

OpenStudy (anonymous):

@swimgirly95 : what is the denominator ?